The function f(x) = 0.0037x2 – 0.46x +36.63 models the average age at which Kentucky men were first married x years after 1900. In which year was this average age at a minimum? What was the average age at first marriage for that year?
Please help me.

Answer by stanbon(75887)

The function f(x) = 0.0037x^2 – 0.46x +36.63 models the average age at which Kentucky men were first married x years after 1900. In which year was this average age at a minimum? What was the average age at first marriage for that year?
-----------
Minimum occurs when x=-b/2a = -(-0.46)/(2(0.0037)) = 62.16
f(62.16216...) = 22.333
Average age was 22 1/3 yrs.
=============================
Cheers,
Stan H.